A credit card company wants to estimate the average balance of its customers to within $15 of the true value, with a 95% confidence. Find the minimum required sample size, if the standard deviation is $43.
Answer #1

Solution:

Given in the question

Margin of error = 15/2 =7.5

Standard deviation = 43

Zalpha/2 = 1.96

Margin of error can be calculated as

Margin of error = Zalpha/2 * SD/sqrt(n)

7.5 = 1.96*43/sqrt(n)

N = (1.96*43/7.5)^2

No. Of sample = (11.2373)^2 = 127